# VisLTR: Visualization-in-the-Loop Table Reasoning

### v1

@@ 6 Evaluation
We conduct quantitative evaluations and case studies for a thorough evaluation of VisLTR. The quantitative evaluation experiments evaluate the alignment ability of our fine-tuned LVLM (Sect. 6.1). Next, we present two cases on actual application scenarios (Sects. 6.2 & 6.3), representing a possible pathway through multi-modal interactions to complete a table reasoning task.
6.1 Effectiveness of Multi-modal Visualization Alignment
We evaluate the alignment performance of our multi-modal visualization alignment model from two perspectives: quantitative evaluation on text-chart retrieval, and user recognition on sketch-chart retrieval.
@@ 6.1.1 Text-Chart Retrieval
To rigorously assess the performance of our fine-tuned visualization alignment model in aligning chart-text pairs, we ensure that the evaluation dataset includes chart types of bar, line, and area chart, as well as all $T_{N}=23$ categories of trend words. Specifically, we randomly select 20 percent of the data for each chart type and each trend category from the chart-text pairing dataset for evaluation.
Metrics. Considering the imbalance among different trend word categories, we set the following metrics for a comprehensive evaluation:
  • Weighted Recall (WR). WR gauges the ability to correctly identify all relevant results among all actual positives. $WR=\frac{1}{T_{N}}\sum_{i=1}^{T_{N}}R_{i}$, where $R_{i}$ is the recall of the $i^{th}$ trend-word category.
  • Weighted Precision (WP). WP determines the proportion of true positives within all positively classified results. $WP=\frac{1}{T_{N}}\sum_{i=1}^{T_{N}}P_{i}$, where $P_{i}$ is the precision of the $i^{th}$ trend-word category.
  • Weighted F1-Score (WF). WF adopts the weighted average considering the unevenness of the categories, which is calculated using the formula: $WF=\frac{1}{T_{N}}\sum_{i=1}^{T_{N}}\frac{2\times P_{i}\times R_{i}}{(P_{i}+R_{i})}.$
@@
Results. The quantitative results are presented in Table 1. The overall results for all chart types confirm the robustness of our model in aligning chart and text modalities. Moreover, the results for each individual chart type also provide substantial evidence of the model’s effectiveness across various chart types. Our model has the effective accuracy, with performance metrics exceeding 85% across all chart types. Nevertheless, it is observed that bar charts present slightly lower metrics in comparison to other chart types. This variation may be attributed to the inherent characteristics of bar charts, which utilize separated bars to encode data, resulting in less distinctiveness compared to the line marks used in line charts and the area marks used in area charts. To enhance the accuracy for bar charts, factors such as bar widths need to be taken into consideration.
@@ 6.1.2 Sketch-Chart Retrieval
We validate the alignment for chart-sketch pairs through a user study. We selected three pre-trained models, CLIP[41], OpenCLIP[8], and UniChart[37] for comparison. Initially, we randomly selected 200 sketches from the chart-sketch dataset. Then, we tasked each of the four models with determining the closest chart in the intent space to each selected sketch, indicating the chart with the highest similarity. This process constructed 800 matched sketch-chart pairings for subsequent user study. We recruited 30 participants (15 women, and 15 men) with experience in building charts. Each participant is randomly given 20 sketches with their matched charts by each morel and asked to rate the similarity between pairs of matched sketches and charts for each model, based on their visual perception of the correspondences. The ratings are gathered using a 5-point Likert scale[22], which provides a quantitative measure of the user agreement or disagreement with the similarities of chart-sketch pairs.
@@
Figure 6: Ratings of the visual similarity between pairs of the matched sketches and charts across our fine-tuned LVLM and three pre-trained models based on a 5-point Likert scale.
@@ 6.2 Case Study 1: Financial Data Exploration
In this case, we consider a financial data exploration scenario. A market analyst, Mike, is studying the Dow Jones Index (DJI.)’s change patterns post-COVID-19. The used data measures the daily price movements of 30 large US companies, spanning from January 2nd, 2020 to December 29th, 2023. The data includes five columns: Date, Open, High, Low, and Close, detailing the trading metrics for each business day. Mike first loads the DJI. data into VisLTR and selects ‘Close’ as the displayed variable in the Main View. VisLTR initially generates 64,261 visualization references, among which 13,823 visualizations are pruned for exploration. The average retrieval time for querying a visualization reference stored in Chroma is within 0.6s.
@@
  • Table Summarization. Mike first enters in the dialogue box “Describe the data.”. The response from VisLTR includes two parts, one is the basic table description “The table shows the market performance of the Dow Jones Index from January 2, 2020 to December 29, 2023, including the date, opening index, intraday high index, intraday low index, and closing index.”, and the other is the overall trend description “The data suffers a sharp drop and then a gradual recovery. There are lots of fluctuations during the overall upward trend.”.
  • Question & Answering. Given this description, Mike is interested in the “sharp drop” pattern, so he tries to locate the pattern by entering the text “Please give me more details about the mentioned ‘sharp drop’ pattern.”. Based on the context, VisLTR can easily locate the “sharp drop” pattern appears, and the matched interval is highlighted in Main View. At the same time, a text description “In March 2020, the Dow Jones Index experienced a ‘sharp drop’ pattern. This drop reflects significant market volatility amid growing concerns over the COVID-19 pandemic.” is returned in the dialogue box.
  • Visual Query. Mike then freely sketches a ‘w’ pattern [Uncaptioned image] in the empty canvas and enters “Are there patterns in the data similar to my hand-drawn sketch?”. VisLTR retrievals for the visualization references similar to the input hand-drawn sketch from the vector database, and returns the three most similar matches in Main View. To facilitate user understanding, VisLTR also returns a text description of the matched results: “The three intervals with similar patterns are from 8-Feb-21 to 16-Mar-21, 18-Feb-21 to 18-Mar-21, and 15-Dec-20 to 11-Jan-21. This pattern is recognized as a double-bottom pattern. It occurred in December 2020 and February 2021 suggesting that during both periods, the market demonstrated buyer interest at lower price levels after experiencing declines. This support prevented further downward movement. Higher prices prompted more buyers to participate, contributing to the Dow’s upward momentum.”.
In this way, Mike gets an overall understanding of the tabular data, and also typical trends and patterns within the data. The multi-modal interactions, together with visual interactions like selecting and zooming in the visual interface, enable Mike to easily understand the DJI. data and identify interesting data change patterns.
Figure 7: Case study 1: data exploration for DJI. form 2020 to 2023. Mike gets a complete understanding of the data via table summarization, question & answering, and visual query enabled by VisLTR.
Figure 8: Case study 2: multivariate analysis for the uploaded air pollutant dataset. With VisLTR, Emily identifies typical data change patterns and relationships between different air pollutants.
@@ 6.3 Case Study 2: Air Pollutant Analysis
In the second case, we demonstrate the workflow of mining relationships between data change patterns of different variables. A student, Emily, is seeking the correlations between different air pollutants. The used air pollutant dataset includes the daily air quality metrics from January 1, 2017 to December 31, 2018, which is collected by the Beijing US Embassy [^1] https://aqicn.org/city/beijing/us-embassy/cn/. The data includes variables of Date, PM25, O3, temperature, and relative humidity, which provides insights into the fluctuations in air quality and correlations between air pollutants. Emily first loads the pollutant data into VisLTR, which generates a multi-line chart with each line representing a variable. In the backend, VisLTR generates a total of 102,787 visualization references, of which 40,243 are stored in Chroma after pruning.
Figure 8 illustrates Emily’s analysis process using VisLTR to explore the multivariate data.
  • Free Exploration. To obtain the typical data change patterns of pollutants, Emily selects four pollutant indicators, PM25, O3, temperature, and relative humidity, and obtains their typical change patterns in the Pattern View. She notices that all the four indicators exhibit obvious daily fluctuations. Specifically, PM25 exhibits a ‘zigzag pattern with varying amplitudes.’, indicative of its dynamic nature and daily variability. Similarly, a typical pattern of O3 is ‘An irregular zigzag pattern with an overall fluctuating trend.’, further highlighting the pronounced daily jitters in pollutant levels. These observations motivate Emily to explore the relationship between these pollutant indicators over a long time interval.
@@
In this manner, Emily gains a comprehensive insight into the multivariate air pollutant data, identifying both the typical data change patterns and correlations among various air pollutant indicators.
@@ 7 Discussion
The case studies have demonstrated VisLTR’s capability in identifying and explaining data change patterns within tabular data. By integrating visualization-in-the-loop, VisLTR not only complements SQL-based LLMs for table reasoning but also aligns with the exploratory data analysis (EDA) process. EDA often involves summarizing the main characteristics of a dataset using visualization, enabling users to understand the data’s structure, identify patterns, detect anomalies, and formulate hypotheses for further investigation. In this sense, the proposed framework not only streamlines the process of EDA but also expands its horizons, providing users with novel approaches to uncover hidden patterns within datasets.
@@
Generalizability. Nevertheless, to fully realize the potential of the proposed framework, we need to enhance the generalizability of VisLTR from the following perspectives.
@@
  • More table types. VisLTR is currently instantiated on time-series tables, which contain data change patterns that cannot be processed by existing SQL-based LLMs for table reasoning. VisLTR has the potential to be extended to more general table types by employing alternative data transformation methods within the visualization referencing module. For instance, VisLTR can be adapted to recognize and analyze the distribution patterns in a student grade table. In this scenario, a data aggregation method could be employed to identify grade distribution patterns, such as the normal distribution. Supporting a broader range of table types is feasible, leveraging methods such as subset embedding[54] and learning-based approaches[46].
@@
  • General user queries. VisLTR focuses on tasks related to reasoning about data change patterns in tables. Therefore, we primarily provide text-based user queries related to data change patterns, such as “what type of change pattern exists in the data?” However, this narrow focus also limits our ability to handle other general user queries, including those related to causal reasoning, such as “why did Apple stock rise in March?” When posed with such questions, VisLTR may encounter the ‘hallucination’ issue, leading to incorrect queries and responses. To achieve a comprehensive exploration and reasoning of a data table, we require more general user queries, which may necessitate the incorporation of external knowledge sources.
@@
  • Intent-enhanced sketches. Users’ intuition about sketching targets often emerges from the exploration process. As a reason, VisLTR offers insights into common patterns within the uploaded table. However, expressing complex and less common patterns solely through hand-drawn sketches can be challenging for users[27]. To address this challenge, several solutions can be considered. First, incorporating additional interactive features such as dragging, resizing, and erasing, sketches can provide users with greater flexibility in adjusting their intent[30]. Summarizing multi-attribute temporal patterns and introducing a more effective time-series segmentation method have shown promise in addressing this issue[44]. By implementing these methods, we can further enhance the capabilities of VisLTR to support sketches with enhanced intents.
@@
Scaling up the dataset. The successful alignment of our model heavily relies on acquiring high-quality training data. Scaling up the dataset to enhance the generalizability of VisLTR presents significant challenges for the data augmentation and user labeling strategies currently employed. Collecting user-labeled sketches is time-consuming, and incorporating intents into sketches would further increase complexity. To address this limitation, one solution is to leverage crowdsourcing strategies, although this still requires significant manual effort. Another innovative technique would be to use large models to simulate and augment a variety of hand-drawn patterns. LLM-powered data augmentation has been a promising direction[9], and we are interested in using this approach for scaling up the dataset.
@@
Potential scenarios. Besides table reasoning, the proposed framework holds potential for other application scenarios, such as safeguarding sensitive data that requires a balance between data privacy and utility. By transforming raw tabular data into visualizations, VisLTR shields the secure information contained within tables from users, while still allowing for examination and querying of the data patterns. Moreover, VisLTR could also find applications in fraud detection, as visualizations generated from transactional data could help identify suspicious patterns indicative of fraudulent activities.
@@ 8 Conclusion
This work introduces VisLTR system built upon a novel visualization-in-the-loop framework for table reasoning, serving as a complement to existing SQL-based table reasoning research. By utilizing visualizations as a proxy, the framework offers several advantages, including the ability to identify data change patterns, mitigate LLM drift, and support multi-modal interactions. The implementation of VisLTR requires an effective multi-modal visualization alignment model that binds chart, text, and sketch modalities together in a joint embedding space. This is accomplished by fine-tuning an LVLM with a new dataset of chart-text and chart-sketch pairings. Together with strategies employed for visualization referencing, indexing, and retrieval, we have created an interactive system that facilitates multi-modal interactions with tabular data. The feasibility of VisLTR is demonstrated through quantitative evaluations of the visualization alignment model and practical case studies. The proposed visualization-in-the-loop framework offers a new research direction for table reasoning. We aim to extend the support to a wider range of table types, accommodate more generalized user queries, and explore additional application scenarios in the future.

### v2

@@ 6 Evaluation
To thoroughly evaluate VisTR, we conduct both quantitative evaluations and case studies. In the quantitative evaluations (Sect. 6.1), we assess the alignment ability of our fine-tuned MLLM. Subsequently, we present two cases on actual application scenarios (Sects. 6.2 & 6.3), representing a possible pathway through multimodal interactions to complete a time-series table reasoning task.
6.1 Effectiveness of Multimodal Visualization Alignment
We evaluate the alignment performance of our multimodal visualization alignment model from two perspectives: quantitative evaluation on text-chart retrieval, and user recognition on sketch-chart retrieval.
6.1.1 Baselines and Metrics
We compare our fine-tuned MLLM against several pre-trained models: CLIP[48], OpenCLIP[10], and UniChart[43]. Our MLLM is a fine-tuned CLIP model on the composite dataset. To highlight the necessity and impact of the fine-tuning process, we compare it with the zero-shot CLIP and OpenCLIP models as an ablation study. Additionally, we validate the effectiveness of our fine-tuned CLIP by comparing it against OpenCLIP and Unichart, confirming the appropriateness of our model selection. VisTR enables comprehensive multimodal binding among different modalities, rather than performing direct subsequence or sketch matching. In this sense, we omit the comparison with existing sketch query[41, 17] or sequence matching systems[21, 67].
Considering the imbalance among different trend word categories, we set two metrics for a comprehensive evaluation:
  • Weighted F1-Score (WF). WF adopts the weighted average considering the unevenness of the categories, which is calculated as: $WF=\frac{1}{T_{N}}\sum_{i=1}^{T_{N}}\frac{2\times P_{i}\times R_{i}}{(P_{i}+R_{i})}$, where $R_{i}$ and $P_{i}$ are the recall and precision of the $i^{th}$ trend-word category, respectively.
@@ 6.1.2 Text-Chart Retrieval
To rigorously assess the performance of our fine-tuned visualization alignment model in aligning chart-text pairs, we ensure that the evaluation dataset includes chart types of bar, line, and area chart, as well as all $T_{N}=23$ categories of trend words. Specifically, we randomly select 20% of the data for each chart type and each trend category from the chart-text pairing dataset for evaluation.
@@
Fine-tuned OpenCLIP Acc (%) 50.74 50.36 51.86 50.24
WF (%) 49.77 48.04 46.41 49.73
Zero-shot CLIP Acc (%) 0.82 1.10 0.72 1.02
WF (%) 0.36 0.29 0.11 0.07
Zero-shot OpenCLIP Acc (%) 7.71 8.20 5.27 1.08
WF (%) 5.37 3.78 1.95 0.30
Results. The quantitative results are presented in Table 1. Our fine-tuned MLLM demonstrates robust performance in aligning chart and text modalities, with overall accuracy and weighted F1-score exceeding 85% across all chart types. The results for each individual chart type also provide substantial evidence of the model’s effectiveness across various chart types. However, we observed that bar charts yield slightly lower metrics than other chart types. This discrepancy may be attributed to the inherent characteristics of bar charts, which utilize separated bars to encode data, resulting in less distinctiveness compared to the continuous line marks used in line charts and the filled area marks used in area charts. To enhance the accuracy of bar charts, factors such as bar widths may need to be considered during fine-tuning process.
The comparisons between fine-tuned CLIP and zero-shot CLIP models confirm that fine-tuning is crucial for enhancing model performance in our specific domain. While zero-shot OpenCLIP slightly outperforms zero-shot CLIP, fine-tuned CLIP significantly surpasses fine-tuned OpenCLIP. Zero-shot OpenCLIP outperforms zero-short CLIP is probably due to more extensive pre-training data that include more diverse chart images. However, after fine-tuning, CLIP’s architecture process is better suited for our specific requirements. Its superior results validate our choice of CLIP as the base model.
Figure 7: Ratings of the visual similarity between pairs of the matched sketches and charts across our fine-tuned MLLM and three pre-trained models based on a 5-point Likert scale.
@@ 6.1.3 Sketch-Chart Retrieval
We validate the alignment for chart-sketch pairs through a user study. Initially, we randomly selected 200 sketches from the chart-sketch dataset. Then, we tasked each of the four models (ours, CLIP, OpenCLIP, and UniChart) with determining the closest chart in the intent space to each selected sketch, indicating the chart with the highest similarity. This process constructed 800 matched sketch-chart pairings for the subsequent user study. We recruited 30 participants (15 women, and 15 men) with experience in building charts. Each participant is randomly given 20 sketches with their matched charts by each morel and asked to rate the similarity between pairs of matched sketches and charts for each model, based on their visual perception of the correspondences. The ratings are gathered using a 5-point Likert scale[27], which provides a quantitative measure of the user agreement or disagreement with the similarities of chart-sketch pairs.
@@ 6.2 Case Study 1: Financial Data Exploration
In this case, we consider a financial data exploration scenario. A market analyst, Mike, is studying the Dow Jones Index (DJI.)’s change patterns post-COVID-19. The used data measures the daily price movements of 30 large US companies, spanning from January 2nd, 2020 to December 29th, 2023. The data includes five columns: Date, Open, High, Low, and Close, detailing the trading metrics for each business day. Mike first loads the DJI. data into VisTR and selects ‘Close’ as the displayed variable in the Main View. VisTR initially generates 64,261 visualization references, among which 13,823 visualizations are pruned for exploration. The average retrieval time for querying a visualization reference stored in Chroma is within 0.6s.
@@
  • Table Summarization. Mike first enters the dialogue box “Describe the data.”. The response from VisTR includes two parts, one is the basic table description “The table shows the market performance of the Dow Jones Index from January 2, 2020, to December 29, 2023, including the date, opening index, intraday high index, intraday low index, and closing index.”, and the other is the overall trend description “The data suffers a sharp drop and then a gradual recovery. There are lots of fluctuations during the overall upward trend.”.
  • Question & Answering. Given this description, Mike is interested in the “sharp drop” pattern, so he tries to locate the pattern by entering the text “Please give me more details about the mentioned ‘sharp drop’ pattern.”. Based on the context, VisTR can easily locate the “sharp drop” pattern that appears, and the matched interval is highlighted in Main View. At the same time, a text description “In March 2020, the Dow Jones Index experienced a ‘sharp drop’ pattern. This drop reflects significant market volatility amid growing concerns over the COVID-19 pandemic.” is returned in the dialogue box.
  • Visual Query. Mike then freely sketches a ‘w’ pattern [Uncaptioned image] in the empty canvas and enters “Are there patterns in the data similar to my hand-drawn sketch?”. VisTR retrievals for the visualization references similar to the input hand-drawn sketch from the vector database, and returns the three most similar matches in Main View. To facilitate user understanding, VisTR also returns a text description of the matched results: “The three intervals with similar patterns are from 8-Feb-21 to 16-Mar-21, 18-Feb-21 to 18-Mar-21, and 15-Dec-20 to 11-Jan-21. This pattern is recognized as a double-bottom pattern. It occurred in December 2020 and February 2021 suggesting that during both periods, the market demonstrated buyer interest at lower price levels after experiencing declines. This support prevented further downward movement. Higher prices prompted more buyers to participate, contributing to the Dow’s upward momentum.”.
In this way, Mike gets an overall understanding of the tabular data, and also typical trends and patterns within the data. The multimodal interactions, together with visual interactions like selecting and zooming in the visual interface, enable Mike to easily understand the DJI. data and recognize interesting data change patterns.
Figure 8: Case study 1: data exploration for DJI. form 2020 to 2023. Mike gets a complete understanding of the data via table summarization, question & answering, and visual query enabled by VisTR.
Figure 9: Case study 2: multivariate analysis for the uploaded air pollutant dataset. With VisTR, Emily recognizes typical data change patterns and relationships between different air pollutants.
@@ 6.3 Case Study 2: Air Pollutant Analysis
In the second case, we demonstrate the workflow of mining relationships between data change patterns of different variables. A student, Emily, is seeking the correlations between different air pollutants. The used air pollutant dataset includes the daily air quality metrics from January 1, 2017, to December 31, 2018, which is collected by the Beijing US Embassy [^1] https://aqicn.org/city/beijing/us-embassy/cn/. The data includes variables of Date, PM25, O3, temperature, and relative humidity, which provides insights into the fluctuations in air quality and correlations between air pollutants. Emily first loads the pollutant data into VisTR, which generates a multi-line chart with each line representing a variable. In the backend, VisTR generates a total of 102,787 visualization references, of which 40,243 are stored in Chroma after pruning.
Figure 9 illustrates Emily’s analysis process using VisTR to explore the multivariate data.
  • Free Exploration. To obtain the typical data change patterns of pollutants, Emily selects four pollutant indicators, PM25, O3, temperature, and relative humidity, and obtains the change patterns in the Pattern View. She notices that all four indicators exhibit obvious daily fluctuations. Specifically, PM25 exhibits a ‘zigzag pattern with varying amplitudes.’, indicative of its dynamic nature and daily variability. Similarly, a typical pattern of O3 is ‘An irregular zigzag pattern with an overall fluctuating trend.’, further highlighting the pronounced daily jitters in pollutant levels. These observations motivate Emily to explore the relationship between these pollutant indicators over a long time interval.
@@
In this manner, Emily gains a comprehensive insight into the multivariate air pollutant data, recognizing both the typical data change patterns and correlations among various air pollutant indicators.
@@ 7 Discussion
The case studies have demonstrated VisTR’s capability in recognizing and explaining data change patterns within tabular data. By leveraging visualizations as representation, VisTR successfully addresses the limitations of existing LLM-based table reasoning methods in pattern recognition and visual-based pattern exploration. Moreover, the process also aligns with the exploratory data analysis (EDA) process. EDA often involves summarizing the main characteristics of a dataset using visualization, enabling users to understand the data’s structure, recognize patterns, detect anomalies, and formulate hypotheses for further investigation. In this sense, the proposed framework not only streamlines the process of EDA but also expands its horizons, providing users with novel approaches to uncover hidden patterns within datasets.
@@
Generalizability. Nevertheless, to fully realize the potential of the proposed framework, we need to enhance the generalizability of VisTR from the following perspectives.
@@
  • More table types. VisTR is currently instantiated on time-series tables, which contain data change patterns that cannot be processed by existing SQL-based LLMs for table reasoning. VisTR has the potential to be extended to more general table types by employing alternative data transformation methods within the visualization referencing module. For instance, VisTR can be adapted to recognize and analyze the distribution patterns in a student grade table. In this scenario, a data aggregation method could be employed to recognize grade distribution patterns, such as the normal distribution. Supporting a broader range of table types is feasible, leveraging methods such as subset embedding[62] and learning-based approaches[54].
@@
  • General user queries. VisTR focuses on tasks related to reasoning about data change patterns in tables. Therefore, we primarily provide text-based user queries related to data change patterns, such as “what type of change pattern exists in the data?” However, this narrow focus also limits our ability to handle other general user queries, including those related to causal reasoning, such as “why did Apple stock rise in March?” When posed with such questions, VisTR may encounter the ‘hallucination’ issue, leading to incorrect queries and responses. To achieve a comprehensive exploration and reasoning of a data table, we require more general user queries, which may necessitate the incorporation of external knowledge sources.
@@
  • Intent-enhanced sketches. Users’ intuition about sketching targets often emerges from the exploration process. As a reason, VisTR offers insights into common patterns within the uploaded table. However, expressing complex and less common patterns solely through hand-drawn sketches can be challenging for users[32]. To address this challenge, several solutions can be considered. First, incorporating additional interactive features such as dragging, resizing, and erasing, sketches can provide users with greater flexibility in adjusting their intent[36]. Summarizing multi-attribute temporal patterns and introducing a more effective time-series segmentation method have shown promise in addressing this issue[51]. By implementing these methods, we can further enhance the capabilities of VisTR to support sketches with enhanced intents.
@@
Scaling up the dataset. The successful alignment of our model heavily relies on acquiring high-quality training data. Scaling up the dataset to enhance the generalizability of VisTR presents significant challenges for the data augmentation and user labeling strategies currently employed. Collecting user-labeled sketches is time-consuming, and incorporating intents into sketches would further increase complexity. To address this limitation, one solution is to leverage crowdsourcing strategies, although this still requires significant manual effort. Another innovative technique would be to use large models to simulate and augment a variety of hand-drawn patterns. LLM-powered data augmentation has been a promising direction[11], and we are interested in using this approach for scaling up the dataset.
@@
Potential scenarios. Besides table reasoning, the proposed framework holds potential for other application scenarios, such as safeguarding sensitive data that requires a balance between data privacy and utility. By transforming raw tabular data into visualizations, VisTR shields the secure information contained within tables from users, while still allowing for examination and querying of the data patterns. Moreover, VisTR could also find applications in fraud detection, as visualizations generated from transactional data could help recognize suspicious patterns indicative of fraudulent activities.
@@ 8 Conclusion
This work introduces VisTR system, built upon a novel framework for table reasoning, expanding upon existing LLM-based table reasoning research. By leveraging visualizations as representations, the framework offers several advantages, including the ability to recognize data change patterns, mitigate LLM drift, and support visual-based exploration. The implementation of VisTR requires an effective multimodal visualization alignment model that binds chart, text, and sketch modalities together in a joint embedding space. This is accomplished by fine-tuning an MLLM with a new dataset of chart-text and chart-sketch pairings. Combined with strategies employed for visualization referencing, indexing, and pruning, we have developed an interactive system that facilitates multimodal interactions with time-series tables. The feasibility of VisTR is demonstrated through quantitative evaluations of the visualization alignment model and practical case studies on time-series data tables. This proposed framework offers a new research direction for table reasoning and enhances the fuzzy exploration process of existing LLM-based table reasoning methods. We aim to extend its support to a wider range of table types, accommodate more generalized user queries, and explore additional application scenarios in the future.